OpenStudy (anonymous):

Select all of the values that are solutions of x2 + 20 = 12x. 1.) x = -10 2.) x = -2 3.) x = 2 4.) x = 10

OpenStudy (anonymous):

just tell me which number choices are right if you would <3

OpenStudy (anonymous):

Plug -10,-2,2,10 in for x. 1) 2(-10)+20 = 12(-10) -20+20=-120 0 = -120 no 2) 2(-2)+20=12(-2) -4+20=-24 16=-24 no 3) 2(2)+20=12(2) 4+20=24 24=24 yes 4) 2(10)+20=12(10) 20+20=120 40=120 no Answer is x=2 Hope that helps

OpenStudy (anonymous):

just x=2 ?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

Was it x^2? or just x*2? in your question

OpenStudy (anonymous):

I may have confused my self with the notation of the equation. Is it x^2 + 20 =12x? or just 2x + 20 = 12x If so, then the answer is x = 2 and x = 10 Sorry for the confusion